OpenStudy (anonymous):

A monkey uses a fishing pole to raise 750g of bananas up to his position 15m above the ground. If the monkey's power output was 330 W, how long did it take to reel in the bananas?

OpenStudy (anonymous):

W=f*d?

OpenStudy (shane_b):

Let me restate that...calculate the PE :)

OpenStudy (anonymous):

Oh, okay. \[PE=mg \Delta y\]?

OpenStudy (shane_b):

Yes \[\Delta PE = mg\Delta h\]

OpenStudy (anonymous):

Okay: \[(750g)(9.81m/s^2)(15m)=110,362.5J\]

OpenStudy (shane_b):

Almost...your mass needs to be in kg. Anyway, you'll get 110.3625J

OpenStudy (anonymous):

So the mass would be 7.5 kg?

OpenStudy (shane_b):

No :) 750g = 0.75kg

OpenStudy (anonymous):

Sorry, what is s?

OpenStudy (shane_b):

seconds

OpenStudy (shane_b):

\[Watts=\frac{Joules}{second}\]

OpenStudy (anonymous):

and that would be equal to 330W

OpenStudy (shane_b):

To be more clear:\[330W=\frac{330Joules}{second}\]

OpenStudy (anonymous):

Hm. That can't be right, can it? Perhaps: \[330W=110.3625J/time (s)\]

OpenStudy (shane_b):

That's not what I'm saying :) I just showed the conversion of 330W to Joules/second

OpenStudy (anonymous):

Oh right, I guess I knew that.

OpenStudy (shane_b):

So if you have 330J/s of power available and need to perform 110.3625J of work, how long would that take?

OpenStudy (anonymous):

36419.63 seconds?

OpenStudy (anonymous):

That seems really long.

OpenStudy (anonymous):

Also seems wrong.

OpenStudy (shane_b):

It is. So your intuition is correct...that's not the answer

OpenStudy (shane_b):

\[\frac{110.3625J}{330J/s}=?\]

OpenStudy (shane_b):

It will take about 1/3 of a second :)

OpenStudy (anonymous):

I got .3344 seconds. Don't know how I got so confused the first time around. Thanks for your help, Shane!

OpenStudy (shane_b):

no problem. I get confused sometimes too...it's physics after all!